I am having difficulty with this problem:
A gift of $6000 grew into $6,000,000 in 200 years. What is the interest rate if it is compounded annually?

Answer by stanbon(75887)

A gift of $6000 grew into $6,000,000 in 200 years. What is the interest rate if it is compounded annually?
--------
A(t) = P(1+(r/n))^(nt)
------
6 million = 6 thousand * (1+r)^(200)
-------
1000 = (1+r)^200
----
200*log(1+r) = log(1000)
------
log(1+r) = 3/200
-----
1+r = 10^(3/200)
1+r = 1.035..
rate = 0.035 = 3.5%
==========================
Cheers,
Stan H.
==========================